1.the classroom is 20 feet long and 30 feet wide.the principal decided that the tiles would look attractive in that class.if each tile is 24 inches long and 36 inches wide,how many tiles are needed to fill the classroom?
2.a rectangle is 4 times as long as it is broad.if the lenth is increased by 4 inches and the width is decreased by 1 inch,the area would be 60 squares inches.what were the dimensions of the original rectangle?

Answer by mananth(16946)

1.the classroom is 20 feet long and 30 feet wide.the principal decided that the tiles would look attractive in that class.if each tile is 24 inches long and 36 inches wide,how many tiles are needed to fill the classroom?
Area of floor = L * W 30 fT820 ft = 600 ft^2
Area of tile = 24 inch * 36 in.
= 3ft*2ft
=6ft2
Number of tiles = area of floor/ area of tile
= 600/6
You continue

2.a rectangle is 4 times as long as it is broad.if the lenth is increased by 4 inches and the width is decreased by 1 inch,the area would be 60 squares inches.what were the dimensions of the original rectangle?
a rectangle is 4 times as long as it is broad.
let breadth be x
length will be 4x
if the length is increased by 4 inches(4x+4) and the width is decreased by 1 inch,the (x-1)area would be 60 squares inches.
(x-1)(4x+4)=60
4x^2+4x-4x-4=60
4x^2=64
x^2=16
x= 4
Breadth = 4 in
length = 16 in